3600 dollars is placed in an account with an annual interest rate of 5.5%. To the

nearest year, how long will it take for the account value to reach 7800 dollars?

1


1
Expert's answer
2021-12-16T15:31:26-0500

Solution:


Answer: 14.44

Step-by-step explanation: The formula says



"A=p(1+r)^t"


A = 7800

P = 3600

R = 0.055

T = ?


"7800=3600(1+0.055)^t"

"\frac{7800}{3600}=(1.055)^t"


"log(\frac{7800}{3600})=t\times log(1.055)"


"T =\frac{ log(\frac{7800}{3600})}{log(1.055)}"


"T = 14.44years"



Answer: 14.44 years